# Disentangled Dynamic Intrusion Detection

### v2

@@ IV-D Classifier and Loss Function
$\mathcal{L}_{\text{Smooth}}=\sum_{t=0}^{T}\left\|\mathbf{X}_{t+\Delta{t}}-\mathbf{X}_{t}\right\|_{2}.$ (19)
Finally, the overall loss of DIDS can be expressed as follows:
$\mathcal{L}=\mathcal{L}_{\text{Int}}+\delta\mathcal{L}_{\text{Smooth }}+\zeta\mathcal{L}_{\text{Dis}},$ (20)
where $\delta$, and $\zeta$ are trade-off parameters.
IV-E Multi-scale Few-shot Learning
So far we have proposed DIDS for supervised intrusion detection. Furthermore, we propose the following MFL to detect few-shot threats accurately, e.g., $5$ samples of each attack type. We first denote the learned representations $\mathbf{X}_{t+\Delta t}$ in Eq. 17 as $\mathbf{Z}\in\mathcal{R}^{L*N}$, where $L$ is the length of representations, and $N$ is the few-shot sample number. The few-shot learner aims to generate a coefficient matrix $\mathbf{S}$ as a learned similarity matrix between each sample pair, thus benefiting the distinction of few-shot traffic threats.
We first focus on the original representations $\mathbf{Z}$ to generate a coefficient matrix in the following formulation:
$\min_{\mathbf{H}}=\|\mathbf{Z}-\mathbf{Z}\mathbf{H}\|_{F}^{2}+\lambda_{1}\|\mathbf{H}\|_{F}^{2},s.t.\rm{diag}(\mathbf{H})=0,$ (21)
where $|\cdot|_{F}^{2}$ is the Frobenius matrix norm (F-norm) and $\rm{diag}(\mathbf{H})$ denotes the diagonal entries of $\mathbf{H}$. $\mathbf{H}$ can be directly derived by solving Eq. 21. Specifically, we directly employ Eq. 21 as a loss function and utilize backpropagation and optimizer, e.g., Adam, to solve it.
The above self-expressive optimization only focuses on preserving the original-scale information of $\mathbf{Z}$. The multi-scale information of $\mathbf{Z}$ can also provide necessary information to make few-shot traffic samples distinguishable. To this end, we proposed a transform-based algorithm to generate coefficient matrix $\mathbf{Q}$ by using $\mathbf{Z}$’s multi-scale information. This process can be considered as an ‘augmentation’ of the original representations $\mathbf{Z}$ in the latent space, aiming to capture the invariant attack-specific features across different scales. Finally, we fuse the obtained coefficient matrices $\mathbf{H}$ and $\mathbf{Q}$. We refer to the above model as Multi-scale Few-shot learning (MFL). Equipped with the fusion of generated matrices, MFL can effectively utilize few-shot representations across different scales, i.e., under the original and transformed scale, thus benefiting the distinction of few-shot traffic threats. We have the following principles to achieve MFL and clarify the motivations behind these designs:
  • 1. MFL should encourage the representations under different scales to be close in latent space, thus capturing the attack-specific features and benefiting the distinction of traffic samples in different types, which is motivated by the empirical studies in Fig. 4 (a)(b)
  • 2. MFL should ensure the disentanglement among the elements of the learned representation, thus highlighting the attack-specific information, which is motivated by the empirical studies in Fig. 4 (c)(d)
To achieve the above goals, we first propose a transformation-based learning scheme by projecting the representations $\mathbf{Z}$ to a latent space, aiming to discover the attack-specific invariant features across different scales. Then we introduce a regularization term to disentangle the few-shot representations. Finally, we propose an alternating optimization algorithm to derive $\mathbf{Q}$ in an end-to-end way. The implementation involves the following three steps:
1) Generating Representation Transformation.
We denote $\mathbf{Z}_{o}=\mathbf{Z}$ as the original representations, and $\mathbf{Z}_{t}=\mathcal{G}(\mathbf{Z})$ as the transformed representations, where $\mathcal{G}$ is a transformation operator. Specifically, we introduce a scaling operator $\mathcal{S}$ as follows:
$\mathcal{S}=\operatorname{diag}(s_{1},s_{2},...,s_{L})=\left[\begin{array}[]{ccc}s_{1}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&s_{L}\end{array}\right],$ (22)
where $s_{1},s_{2},...,s_{D}$ is the scale factor, and different values determine different scalars. Then the scaling transformation can be written as $\mathbf{Z}_{t}=\mathcal{G}(\mathbf{Z})=\mathcal{S}\mathbf{Z}$. Especially, when $s_{1}=s_{2}=...=s_{L}=\gamma$, $\mathcal{S}$ is an equal-rate scaling operator with the rate of $\gamma$. Hence, the transformation can be rewritten as $\mathbf{Z}_{t}=\mathcal{G}(\mathbf{Z})=\mathcal{S}\mathbf{Z}=\gamma\mathbf{I}\mathbf{Z}=\gamma\mathbf{Z}$. By the equal-rate transformation, we generate transformed representations while preserving their content, e.g., scaling up an image by the rate of 2, will not change the content we see.
2) Generating Multi-scale Coefficient Matrix.
So far, we have generated the transformations of $\mathbf{Z}$. In step $2$, we propose a constraint term to encourage the representation to be close between the original and transformed scale. Furthermore, we introduce a learnable projection operator $\mathbf{P}$ to uncover the attack-specific invariant information, thus benefiting the distinction of few-shot traffic threats.
$f_{1}(\mathbf{P})=\left\|\mathbf{P}^{T}\mathbf{Z}_{o}-\mathbf{P}^{T}\mathbf{Z}_{t}\right\|_{F}^{2},$ (23)
The projection operator $\mathbf{P}$ can be considered as a metric matrix across different-scale representation, also benefiting the connectivity of same-attack traffic samples under the transformation. Then we formally propose an optimization problem as follows:
$\min f_{1}(\mathbf{P})+\eta\|\mathbf{Q}\|_{F},$ (24)
[…]
The above formulation Eq. LABEL:eq:_Q_2 corresponds to design principle 1 and is motivated by the empirical studies in Fig. 4 (a)(b), aiming to distinguish the attack-specific information in the latent space.
3) Disentangling the Learned Representations.
Finally, we introduce a disentangle regularization term to highlight the attack-specific information, thus mitigating the representation entangled problem in few-shot traffic samples.
$f_{2}(\mathbf{P})=-\sum_{j}\left\|\mathbf{P}^{T}\mathbf{Z}_{o}^{+j}-\mathbf{P}^{T}\mathbf{Z}_{t}^{-j}\right\|_{F}^{2},$ (25)
where $\mathbf{Z}_{o}^{+j}\in\mathbb{R}^{(L-1)\times N}$ repeats the $j$-th row of $\mathbf{Z}_{o}$ and $\mathbf{Z}_{t}^{-j}\in\mathbb{R}^{(L-1)\times N}$ denotes the transformation matrix by removing the $j$-th row. The above disentanglement regularization term Eq. 25 corresponds to design principle 2 and is motivated by the empirical studies in Fig. 4 (c)(d), aiming to highlight the attack-specific information.
Combining the above key formulations Eq. LABEL:eq:_Q_2 and Eq. 25, we can rewrite the optimization problem as:
$\min f(\mathbf{P})+\eta\|\tilde{\mathbf{Q}}\|_{F},$ (26)
[…]
where $f(\mathbf{P})=\frac{\alpha}{2}(f_{1}(\mathbf{P})+f_{2}(\mathbf{P}))$, $\alpha$ and $\eta$ are trade-off hyperparameters, $\tilde{Q}$ is a auxiliary variable, and $\gamma$ is scaling rate hyperparameter.
Since the optimization problem in Eq. LABEL:MSF_2 is not convex with the unknowns $\{\mathbf{P},\mathbf{Q}\}$, we solve $\mathbf{Q}$ by iteratively updating variables while fixing another. We propose the following alternating solution to derive the coefficient matrix $\mathbf{Q}$.
Alternating Solution for MFL
We solve Eq. LABEL:MSF_2 by converting the original problem to the augmented Lagrange minimizing problem, as Eq. LABEL:MSF_2 involving a multi-objective optimization:
$\mathcal{L}=\frac{\alpha}{2}\left\|\mathbf{P}\mathbf{Z}-\gamma\mathbf{P}\mathbf{Z}\mathbf{Q}\right\|_{F}^{2}$ (27)
$-\frac{\alpha}{2}\sum_{j}\left\|\mathbf{P}\mathbf{Z}^{+j}-\gamma\mathbf{P}\mathbf{Z}^{-j}\right\|_{F}^{2}+\eta\|\tilde{\mathbf{Q}}\|_{F}$
$+<\Phi_{1},\mathbf{P}\mathbf{Z}-\gamma\mathbf{P}\mathbf{Z}\mathbf{Q}+<\Phi_{2},\mathbf{Q}-\tilde{\mathbf{Q}}>$
$+\frac{\mu}{2}\left(\left\|\mathbf{P}\mathbf{Z}-\gamma\mathbf{P}\mathbf{Z}\mathbf{Q}\right\|_{F}^{2}+\|\mathbf{Q}-\tilde{\mathbf{Q}}\|_{F}^{2}\right),$
where $\Phi_{1}$ and $\Phi_{2}$ are the Lagrange multipliers and the penalty parameter $\mu>0$.
Then, the alternating solving and updating of each variable can be involved in the following four steps:
1) Update $\mathbf{P}$ by fixing $\mathbf{Q}$ and $\Phi_{1}$.
[…]
The closed-form solution to the problem in Eq. (27) can be given as follows:
$\mathbf{P}^{*}=-\Phi_{1}\Delta_{3}^{T}(\alpha\Delta_{1}\Delta_{1}^{T}-\lambda\sum_{j}\Delta_{2}\Delta_{2}^{T}+\mu\Delta_{3}\Delta_{3}^{T}+\mathbf{I})^{-1}{,}$ (29)
where $\Delta_{1}=(1-\gamma)\mathbf{Z},\Delta_{2}=\mathbf{Z}^{+j}-\mathbf{Z}^{-j},\Delta_{3}=\mathbf{Z}-\gamma\mathbf{Z}\mathbf{Q}$, and $\mathbf{I}$ is an identity matrix with proper size.
2) Update $\mathbf{Q}$ by fixing $\mathbf{P},\tilde{\mathbf{Q}}$ and $\Phi_{1},\Phi_{2}$.
The optimization problem is shown as:
$\mathbf{Q}^{*}$ $=\arg\min_{\mathbf{Q}}\frac{\mu}{2}\left\|\mathbf{P}^{T}\mathbf{Z}-\gamma\mathbf{P}^{T}\mathbf{Z}\mathbf{Q}+\frac{\Phi_{1}}{\mu}\right\|_{F}^{2}$ (30)
$+\frac{\mu}{2}\left\|\mathbf{Q}-\tilde{\mathbf{Q}}+\frac{\Phi_{2}}{\mu}\right\|_{F}^{2}.$
The closed-form solution to the problem above is:
$\mathbf{Q}^{*}=\left(\mu\gamma^{2}\mathbf{P}^{T}\mathbf{ZZ}^{T}\mathbf{P}+\mu\mathbf{I}\right)^{-1}\left(\mu\gamma\mathbf{Z}^{T}\mathbf{P}\Delta_{3}+\mu\tilde{\mathbf{Q}}-\Phi_{2}\right),$ (31)
where $\Delta_{3}=\mathrm{P}^{T}\mathbf{Z}+\frac{\Phi_{1}}{\mu}$.
3) Update the auxiliary variable $\tilde{\mathbf{Q}}$.
It can be solved via the singular value thresholding method,
$\tilde{\mathbf{Q}}^{*}=\arg\min_{\tilde{\mathbf{Q}}}\|\tilde{\mathbf{Q}}\|_{*}+\frac{\mu}{2}\left\|\mathbf{Q}-\tilde{\mathbf{Q}}+\frac{\Phi_{2}}{\mu}\right\|_{F}^{2},$ (32)
and the closed-form solution of $\tilde{\mathbf{Q}}^{*}$ is
$\tilde{\mathbf{Q}}^{*}$ $=\mathcal{S}_{1/\mu}\left(\mathbf{Q}+\Phi_{2}/\mu\right)$ (33)
$=\mathbf{U}\left\{\operatorname{sign}\left(\Sigma_{ii}\right)\max\left(\left|\Sigma_{ii}-1/\mu\right|,0\right)\right\}\mathbf{V}^{T},$
where $\mathcal{S}$ is the singular value thresholding operator, $\operatorname{sign}(\cdot)$ is the indicator function, and $\tilde{\mathbf{Q}}+\Phi_{2}/\mu=\mathbf{U\Sigma}\mathbf{V}^{T}$ is the singular value decomposition of $\tilde{\mathbf{Q}}+\Phi_{2}/\mu$.
4) Update the multipliers $\Phi_{1},\Phi_{2},\mu$.
$\Phi_{1}$ $=\Phi_{1}+\mu\left(\mathbf{P}^{T}\mathbf{Z}-\gamma\mathbf{P}^{T}\mathbf{Z}\mathbf{Q}\right),$ (34)
$\Phi_{2}$ $=\Phi_{2}+\mu(\mathbf{Q}-\tilde{\mathbf{Q}}),$
$\mu$ $=\min\left(1.01\mu,\mu_{\max}\right).$
We can obtain the coefficient matrix $\mathbf{Q}$ following the above updating rules. The pseudo-code of MFL is proposed in Algorithm 1.
0:  $\mathbf{Z}_{o}=\mathbf{Z}$, scaling rate $\gamma$, $\alpha$ and $\eta$. 0:  Affinity matrix $\mathbf{S}$ for later few-shot intrusion detection. 1:  Initialization: $\mathbf{Z}_{t}=\gamma\mathbf{Z}$ by $\gamma$-rate scaling transform on $\mathbf{Z}_{o}$; $\mathbf{Q}=\tilde{\mathbf{Q}}=\mathbf{0}$; and $\Phi_{1}=\Phi_{2}=0;\mu=0.1,\mu_{\max}=10^{7}$. 2:  while Convergence conditions are not satisfied do 3:     Update: $\mathbf{P}$ via Eq. 29. 4:     Update: $\mathbf{Q}$ via Eq. 30. 5:     Update: $\tilde{\mathbf{Q}}$ via Eq. 33. 6:     Update: $\Phi_{1},\Phi_{2}$ and $\mu$ via Eq. 34. 7:     Check the convergence conditions: $\left\|\mathbf{P}\mathbf{Z}-\gamma\mathbf{P}\mathbf{Z}\mathbf{Q}\right\|_{\infty}<10^{-6}\text{. }$ 8:  end while 9:  Obtain the coefficient matrix $\mathbf{Q}$
Algorithm 1 Alternating Algorithm for MFL
So far we have obtained coefficient matrices $\mathbf{H}$ and $\mathbf{Q}$ by parallel computing of Eq. 21 and Algorithm 1. We can fuse two coefficient matrices by introducing a trade-off hyperparameter $\epsilon$:
$\mathbf{Q^{*}}=\mathbf{H}+\epsilon*\mathbf{Q}$ (35)
where $\mathbf{Q}^{*}$ contains fine-grained multi-scale information to help few-shot traffic threats distinguishable.
Finally, we can generate the coefficient matrix of MFL by $\mathbf{S^{*}}=\left(|\mathbf{Q^{*}}|+|\mathbf{Q^{*}}|^{T}\right)/2$. Given the noise information and outliers in the obtained $\mathbf{S}^{*}$, in practical implementations, we further conduct SVD decomposition of $\mathbf{S}^{*}$ to filter noisy information and generate a normalized matrix $\mathbf{S}$. The detailed implementation pseudo-code of generating $\mathbf{S}$ is available in the Appendix $15$.
DIDS-MFL Loss
The DIDS-MFL loss consists of two components, a multi-scale few-shot learning term and a regularization term. The former uses a cross-entropy loss, aiming to match the prediction to the query set via the generated multi-scale coefficient matrix $\mathbf{S}$. In the few-shot intrusion detection task, $\mathbf{S}$ is derived from the representations of support and query samples. The latter enforces the representations $\mathbf{Z}$ to be close to the original-scale representations in latent space. Finally, we introduce a trade-off hyperparameter $\beta$ to control the regularization intensity. The specific DIDS-MFL loss is as follows:
$\mathcal{L}=\sum_{i=1}^{\mathcal{N}}\sum_{j=1}^{\mathcal{Q}}Y_{i,j}\log(\mathrm{P}_{i,j})+\beta\|\mathbf{Z}\mathbf{H}-\mathbf{Z}\|$ (36)
where $\mathcal{N}$ is the category number of attacks, $\mathcal{Q}$ is sample number of query set in each category, $Y_{i,j}$ and $\operatorname{P}_{i,j}$ are ground-truth and prediction, respectively, $\mathrm{P}_{ij}$ is query sample’s mean similarity score derived from the coefficient matrix $\mathbf{S}$, $\beta$ is a trade-off hyperparameter. In Eq. 36, the first term discovers the multi-scale information of $\mathbf{Z}$, while the second term preserves the information under the original scale via the coefficient matrix $\mathbf{H}$. We can fine-tune $\beta$ to control these two terms.
V Experiments
V-A Experimental Settings
  • CIC-ToN-IoT : This dataset is generated from the existing ToN-IoT dataset by a network traffic tool CICFlowMeter, where TON-IoT is a well-known database for intrusion detection collected from Telemetry datasets of IoT services. This dataset consists of $5,351,760$ flows, with $53.00\%$ attack samples and $47.00\%$ benign samples.
  • CIC-BoT-IoT: It is generated from the existing BoT-IoT dataset by CICFlowMeter. This dataset consists of $6,714,300$ traffic flows, with $98.82\%$ attack samples and $1.18\%$ benign samples.
  • EdgeIIoT: It is collected from an IoT/IIoT system that contains mobile devices and sensors. This dataset includes $1,692,555$ flows, with $21.15\%$ attack samples and $78.85\%$ benign samples.
  • NF-UNSW-NB15-v2: It is NetFlow-based and generated from the UNSW-NB15 dataset, which has been expanded with additional NetFlow features and labeled with respective attack categories. This dataset includes $2,390,275$ flows, with $3.98\%$ attack samples and $96.02\%$ benign samples.
  • NF-CSE-CIC-IDS2018-v2: It is a NetFlow-based dataset generated from the original pcap files of CSE-CIC-IDS2018 dataset. This dataset includes $18,893,708$ flows, with $11.95\%$ attack samples and $88.05\%$ benign samples.
Baselines: To evaluate the performance of the proposed DIDS, we select $10$ deep learning based-models as baselines, including $3$ sequence models (i.e., MLP [86], MStream [16], LUCID [37]), $4$ static GCN models (i.e., GAT [87] and E-GraphSAGE [88], SSDCM [89], DMGI [90]), where SSDCM and DMGI are designed for static multi-layer graphs, $4$ dynamic GCN models (i.e., TGN [91], EULER [41], AnomRank [92], DynAnom [92]). Additionally, we choose $3$ rule-based baselines to compare with the proposed DIDS, (i.e., ML [93], AdaBoost [94], and Logistic Regression).
Metrics: We follow the previous works [95] to evaluate the performances of all baselines by two commonly used metrics in intrusion detection including F1-score (F1) and ROC-AUC score (AUC).
TABLE I: Comparisons of binary classification on five datasets. The results with $\ddagger$ are directly copied from [96].
@@ V-A Experimental Settings
TGN [91] $89.90$$\pm$$1.66$ $82.09$$\pm$$1.36$ $96.84$$\pm$$0.44$ $94.41$$\pm$$0.81$ $94.99$$\pm$$0.61$ $89.50$$\pm$$2.04$ $93.55$$\pm$$0.23$ $88.01$$\pm$$1.97$ $95.11$$\pm$$0.46$ $91.30$$\pm$$0.76$
EULER [41] $89.73$$\pm$$1.13$ $80.48$$\pm$$2.46$ $96.00$$\pm$$0.29$ $91.47$$\pm$$1.36$ $92.89$$\pm$$0.32$ $90.64$$\pm$$1.80$ $92.76$$\pm$$0.86$ $86.97$$\pm$$1.11$ $95.87$$\pm$$0.51$ $90.76$$\pm$$0.64$
  AnomRank [92] $76.51$$\pm$$0.98$ $77.41$$\pm$$1.64$ $84.84$$\pm$$0.49$ $82.50$$\pm$$0.59$ $78.37$$\pm$$0.43$ $81.36$$\pm$$0.41$ $90.54$$\pm$$2.40$ $79.63$$\pm$$0.12$ $90.08$$\pm$$0.82$ $83.76$$\pm$$0.35$
DynAnom [97] $79.23$$\pm$$1.81$ $75.22$$\pm$$0.92$ $83.25$$\pm$$0.62$ $79.04$$\pm$$0.84$ $81.56$$\pm$$0.94$ $83.94$$\pm$$0.36$ $89.11$$\pm$$1.48$ $85.25$$\pm$$0.64$ $91.21$$\pm$$0.95$ $88.79$$\pm$$0.54$
Anomal-E [96] - - - - - - 91.8991.89$\ddagger$91.89 ‡ - 94.5194.51$\ddagger$94.51 ‡ -
GAT [87] $86.30$$\pm$$1.16$ $74.66$$\pm$$1.37$ $94.56$$\pm$$0.75$ $93.09$$\pm$$2.83$ $93.30$$\pm$$0.14$ $88.30$$\pm$$1.56$ $92.20$$\pm$$1.60$ $89.91$$\pm$$0.62$ $96.08$$\pm$$0.24$ $90.56$$\pm$$0.34$
E-GraphSAGE [17] $89.46$$\pm$$1.25$ $79.56$$\pm$$1.63$ $93.74$$\pm$$0.76$ $90.53$$\pm$$1.90$ $92.10$$\pm$$1.46$ $89.10$$\pm$$0.64$ $94.10$$\pm$$0.33$ $90.39$$\pm$$0.26$ $95.71$$\pm$$0.35$ $90.22$$\pm$$0.48$
DMGI [90] $88.83$$\pm$$0.48$ $79.13$$\pm$$2.11$ $96.07$$\pm$$1.89$ $92.65$$\pm$$1.57$ $93.83$$\pm$$1.67$ $86.03$$\pm$$2.45$ $93.11$$\pm$$0.98$ $88.51$$\pm$$1.00$ $93.87$$\pm$$0.84$ $87.56$$\pm$$0.55$
SSDCM [89] $89.23$$\pm$$0.87$ $80.84$$\pm$$2.32$ $97.11$$\pm$$0.63$ $94.82$$\pm$$0.96$ $94.72$$\pm$$1.59$ $86.69$$\pm$$0.76$ $93.30$$\pm$$0.25$ $89.22$$\pm$$1.94$ $94.96$$\pm$$0.52$ $88.61$$\pm$$0.38$
MLP [86] $80.74$$\pm$$0.43$ $61.80$$\pm$$1.48$ $93.01$$\pm$$0.60$ $87.90$$\pm$$0.54$ $88.78$$\pm$$0.44$ $86.00$$\pm$$1.49$ $93.12$$\pm$$0.64$ $89.92$$\pm$$0.55$ $94.59$$\pm$$0.94$ $90.42$$\pm$$0.89$
MStream [16] $73.90$$\pm$$1.13$ $70.22$$\pm$$1.61$ $78.48$$\pm$$0.19$ $74.04$$\pm$$1.66$ $82.47$$\pm$$1.67$ $77.89$$\pm$$0.58$ $89.47$$\pm$$1.13$ $84.38$$\pm$$1.01$ $88.34$$\pm$$0.45$ $83.66$$\pm$$1.79$
LUCID [37] $83.62$$\pm$$1.69$ $72.31$$\pm$$1.14$ $94.36$$\pm$$0.41$ $89.46$$\pm$$0.72$ $88.94$$\pm$$1.73$ $85.23$$\pm$$0.94$ $92.77$$\pm$$1.39$ $88.32$$\pm$$0.91$ $95.84$$\pm$$1.46$ $90.75$$\pm$$0.79$
Ours (DIDS) 91.57$\pm$0.40 84.06$\pm$1.01 98.24$\pm$0.32 96.32$\pm$0.25 96.83$\pm$0.36 92.34$\pm$1.10 95.45$\pm$0.67 91.55$\pm$1.03 96.34$\pm$0.21 93.23$\pm$1.50

### v3

@@ 4.4. Classifier and Loss Function
(19) $\mathcal{L}_{\text{Smooth}}=\sum_{t=0}^{T}\left\|\mathbf{X}_{t+\Delta{t}}-\mathbf{X}_{t}\right\|_{2}.$
Finally, the overall loss of 3D-IDS can be expressed as follows:
(20) $\mathcal{L}=\mathcal{L}_{\text{Int}}+\alpha\mathcal{L}_{\text{Smooth }}+\beta\mathcal{L}_{\text{Dis}},$
where $\alpha$, and $\beta$ are trade-off parameters.
Table 1. Comparisons of binary classification on five datasets. The results with $\ddagger$ are directly copied from (Caville et al., 2022).
@@
TGN (Rossi et al., 2020) $89.90$$\pm$$1.66$ $82.09$$\pm$$1.36$ $96.84$$\pm$$0.44$ $94.41$$\pm$$0.81$ $94.99$$\pm$$0.61$ $89.50$$\pm$$2.04$ $93.55$$\pm$$0.23$ $88.01$$\pm$$1.97$ $95.11$$\pm$$0.46$ $91.30$$\pm$$0.76$
EULER (King and Huang, 2022) $89.73$$\pm$$1.13$ $80.48$$\pm$$2.46$ $96.00$$\pm$$0.29$ $91.47$$\pm$$1.36$ $92.89$$\pm$$0.32$ $90.64$$\pm$$1.80$ $92.76$$\pm$$0.86$ $86.97$$\pm$$1.11$ $95.87$$\pm$$0.51$ $90.76$$\pm$$0.64$
AnomRank (Yoon et al., 2019) $76.51$$\pm$$0.98$ $77.41$$\pm$$1.64$ $84.84$$\pm$$0.49$ $82.50$$\pm$$0.59$ $78.37$$\pm$$0.43$ $81.36$$\pm$$0.41$ $90.54$$\pm$$2.40$ $79.63$$\pm$$0.12$ $90.08$$\pm$$0.82$ $83.76$$\pm$$0.35$
DynAnom (Guo et al., 2022) $79.23$$\pm$$1.81$ $75.22$$\pm$$0.92$ $83.25$$\pm$$0.62$ $79.04$$\pm$$0.84$ $81.56$$\pm$$0.94$ $83.94$$\pm$$0.36$ $89.11$$\pm$$1.48$ $85.25$$\pm$$0.64$ $91.21$$\pm$$0.95$ $88.79$$\pm$$0.54$
Anomal-E (Caville et al., 2022) - - - - - - 91.8991.89$\ddagger$91.89 ‡ - 94.5194.51$\ddagger$94.51 ‡ -
GAT (Veličković et al., 2017) $86.30$$\pm$$1.16$ $74.66$$\pm$$1.37$ $94.56$$\pm$$0.75$ $93.09$$\pm$$2.83$ $93.30$$\pm$$0.14$ $88.30$$\pm$$1.56$ $92.20$$\pm$$1.60$ $89.91$$\pm$$0.62$ $96.08$$\pm$$0.24$ $90.56$$\pm$$0.34$
E-GraphSAGE (Lo et al., 2022) $89.46$$\pm$$1.25$ $79.56$$\pm$$1.63$ $93.74$$\pm$$0.76$ $90.53$$\pm$$1.90$ $92.10$$\pm$$1.46$ $89.10$$\pm$$0.64$ $94.10$$\pm$$0.33$ $90.39$$\pm$$0.26$ $95.71$$\pm$$0.35$ $90.22$$\pm$$0.48$
DMGI (Park et al., 2020) $88.83$$\pm$$0.48$ $79.13$$\pm$$2.11$ $96.07$$\pm$$1.89$ $92.65$$\pm$$1.57$ $93.83$$\pm$$1.67$ $86.03$$\pm$$2.45$ $93.11$$\pm$$0.98$ $88.51$$\pm$$1.00$ $93.87$$\pm$$0.84$ $87.56$$\pm$$0.55$
SSDCM (Mitra et al., 2021) $89.23$$\pm$$0.87$ $80.84$$\pm$$2.32$ $97.11$$\pm$$0.63$ $94.82$$\pm$$0.96$ $94.72$$\pm$$1.59$ $86.69$$\pm$$0.76$ $93.30$$\pm$$0.25$ $89.22$$\pm$$1.94$ $94.96$$\pm$$0.52$ $88.61$$\pm$$0.38$
MLP (Roopak et al., 2019) $80.74$$\pm$$0.43$ $61.80$$\pm$$1.48$ $93.01$$\pm$$0.60$ $87.90$$\pm$$0.54$ $88.78$$\pm$$0.44$ $86.00$$\pm$$1.49$ $93.12$$\pm$$0.64$ $89.92$$\pm$$0.55$ $94.59$$\pm$$0.94$ $90.42$$\pm$$0.89$
MStream (Bhatia et al., 2021) $73.90$$\pm$$1.13$ $70.22$$\pm$$1.61$ $78.48$$\pm$$0.19$ $74.04$$\pm$$1.66$ $82.47$$\pm$$1.67$ $77.89$$\pm$$0.58$ $89.47$$\pm$$1.13$ $84.38$$\pm$$1.01$ $88.34$$\pm$$0.45$ $83.66$$\pm$$1.79$
LUCID (Doriguzzi-Corin et al., 2020) $83.62$$\pm$$1.69$ $72.31$$\pm$$1.14$ $94.36$$\pm$$0.41$ $89.46$$\pm$$0.72$ $88.94$$\pm$$1.73$ $85.23$$\pm$$0.94$ $92.77$$\pm$$1.39$ $88.32$$\pm$$0.91$ $95.84$$\pm$$1.46$ $90.75$$\pm$$0.79$
Ours (3D-IDS) 91.57$\pm$0.40 84.06$\pm$1.01 98.24$\pm$0.32 96.32$\pm$0.25 96.83$\pm$0.36 92.34$\pm$1.10 95.45$\pm$0.67 91.55$\pm$1.03 96.34$\pm$0.21 93.23$\pm$1.50
5. Experiments
5.1. Experimental Settings
  • CIC-ToN-IoT (Moustafa, 2019): This dataset is generated from the existing ToN-IoT dataset (Moustafa, 2021) by a network traffic tool CICFlowMeter (Lashkari et al., 2017), where TON-IoT is a well-known database for intrusion detection collected from Telemetry datasets (Alsaedi et al., 2020) of IoT services. This dataset consists of $5,351,760$ flows, with $53.00\%$ attack samples and $47.00\%$ benign samples.
  • CIC-BoT-IoT (Sarhan et al., 2021a): It is generated from the existing BoT-IoT dataset (Koroniotis et al., 2019) by CICFlowMeter. This dataset consists of $6,714,300$ traffic flows, with $98.82\%$ attack samples and $1.18\%$ benign samples.
  • EdgeIIoT (Ferrag et al., 2022): It is collected from an IoT/IIoT system that contains mobile devices and sensors. This dataset includes $1,692,555$ flows, with $21.15\%$ attack samples and $78.85\%$ benign samples.
  • NF-UNSW-NB15-v2 (Sarhan et al., 2022b): It is NetFlow-based and generated from the UNSW-NB15 (Moustafa and Slay, 2015) dataset, which has been expanded with additional NetFlow features and labeled with respective attack categories. This dataset includes $2,390,275$ flows, with $3.98\%$ attack samples and $96.02\%$ benign samples.
  • NF-CSE-CIC-IDS2018-v2 (Sarhan et al., 2022b): It is a NetFlow-based dataset generated from the original pcap files of CSE-CIC-IDS2018 (Sharafaldin et al., 2018) dataset. This dataset includes $18,893,708$ flows, with $11.95\%$ attack samples and $88.05\%$ benign samples.
Baselines: To evaluate the performance of the proposed 3D-IDS, we select $10$ deep learning based-models as baselines, including $3$ sequence models (i.e., MLP (Roopak et al., 2019), MStream (Bhatia et al., 2021), LUCID (Doriguzzi-Corin et al., 2020)), $4$ static GCN models (i.e., GAT (Veličković et al., 2017) and E-GraphSAGE (Hamilton et al., 2017), SSDCM (Mitra et al., 2021), DMGI (Park et al., 2020)), where SSDCM and DMGI are designed for static multi-layer graphs, $4$ dynamic GCN models (i.e., TGN (Rossi et al., 2020), EULER (King and Huang, 2022), AnomRank (Yoon et al., 2019), DynAnom (Yoon et al., 2019)). Additionally, we choose $3$ rule-based baselines to compare with the proposed 3D-IDS, (i.e., ML (Sarhan et al., 2021b), AdaBoost (Lalouani and Younis, 2021), and Logistic Regression).
Metrics: We follow the previous works (Sarhan et al., 2022a) to evaluate the performances of all baselines by two commonly used metrics in intrusion detection including F1-score (F1) and ROC-AUC score (AUC).